The finance manager of ABC ltd has computed the after tax marginal costs of the company;s three sources of capital ordinary shares-600000 at a cost of 14%600000-1500000 at a cost of 17%over 1500000 at cost of 18%preference shares0-300000 at a cost of 12%over 300000 at a cost of 13%debts0-1000000 at a cost of 5%over 1000000 at a cost of 8%the manager wants to maintain the current present capital structure of 50% debt , 10% preference shares and 40% ordinary shares.Requiredi) the break point in the capital structureii) the weighted average cost of capital between breakpoints iii) draw the weighted marginal cost of capital schedule
Answer #1

Following are the given cost of capital of three sources of capital :

  1. Ordinary shares

from 0 to 600000 @ 14%

from 600000 to 1500000 @ 17%

over 1500000 @ 18%

  1. Preference shares

from 0 to 300000 @ 12%

over 300000 @ 13%

  1. Debt

from 0 to 1000000 @ 5%

over 1000000 @ 8%

  1. Capital structure policy ratio

Debt 50%, Preference shares 10% & Ordinary shares 40%

Answer – (i)

Break point in the capital structure shall be the capital structure in which the cost of capital shall be optimum with the given capital structure ratio with the maximum availability and utilization of funds. In the above given cost of capital the highest cost of capital is of ordinary share in which the minimum rate is of 14% limited to a capital of 600000, hence the ordinary share capital would be 600000 to optimize the cost of capital.

To maintain the required capital structure ratio of 10% of total capital to preference shares the preference share capital would be of 150000 at the cost of 12% to optimize the cost of capital.

To maintain the required capital structure ratio of 50% of total capital to debt the debt capital would be of 750000 at the cost of 5% to optimize the cost of capital.

Hence following is the break point capital structure of ABC Ltd.

Ordinary Share Capital

600000

14%

Preference Share Capital

150000

12%

Debt

750000

5%

Answer – (ii)

Weighted average cost of capital between breakpoints shall be calculated using the below mentioned formula :

Weighted average cost of capital = Ke * W1 + Kp * W2 + Kd * W3

Where –

Ke is cost of equity capital – 14%

Kp is cost of preference capital – 12%

Kd is cost of debt – 5%

W1, W2, W3 are the weights calculated in (i) above

Putting the above in formula, we get –

Weighted average cost of capital = Ke * W1 + Kp * W2 + Kd * W3

                                                      = 14 * 6 + 12 * 1.5 + 5 * 7.5

15            15   15

                                                       = 5.6 + 1.2 + 2.5

                                                       = 9.30%

Answer – (iii)

For computing weighted marginal cost of capital schedule we have to assume capital of different amount so as to compute the capital structure and weighted marginal cost of capital in the given scenario

Amount of Capital

Equity Share Capital (40%)

Preference Share Capital (10%)

Debt (50%)

Cost of Equity         (in %)

Cost of Preference (in %)

Cost of Debt     (in %)

Weighted Average Cost of Capital   (in %)

1500000

600000

150000

750000

0.4*14=5.6

0.1*12=1.2

0.5*5=2.5

9.3

2000000

800000

200000

1000000

0.4*17=6.8

0.1*12=1.2

0.5*5=2.5

10.5

2500000

1000000

250000

1250000

0.4*17=6.8

0.1*12=1.2

0.5*5=2.5

10.5

3000000

1200000

300000

1500000

0.4*17=6.8

0.1*12=1.2

0.5*8=4

12